# Augmenting Replay in World Models for Continual Reinforcement Learning

### 30 Jan 2024 v1

@@ Abstract
In continual RL, the environment of a reinforcement learning (RL) agent undergoes change. A successful system should appropriately balance the conflicting requirements of retaining agent performance on already learned tasks, stability, whilst learning new tasks, plasticity. The first-in-first-out buffer is commonly used to enhance learning in such settings but requires significant memory. We explore the application of an augmentation to this buffer which alleviates the memory constraints, and use it with a world model model-based reinforcement learning algorithm, to evaluate its effectiveness in facilitating continual learning. We evaluate the effectiveness of our method in Procgen and Atari RL benchmarks and show that the distribution matching augmentation to the replay-buffer used in the context of latent world models can successfully prevent catastrophic forgetting with significantly reduced computational overhead. Yet, we also find such a solution to not be entirely infallible, and other failure modes such as the opposite — lacking plasticity and being unable to learn a new task — to be a potential limitation in continual learning systems.
@@ 1 Introduction
Over a lifetime, a person or intelligent system that interacts with a complex environment can expect to acquire new experiences and learn new tasks, all while old experiences and tasks are retained. This requirement to satisfy two orthogonal goals — the acquisition of new knowledge and the retention of old knowledge is the challenge of continual learning (CL) Ring (1994), and the demand for both desiderata presents the stability-plasticity dilemma Carpenter and Grossberg (1987). A system capable of successfully undertaking CL should meet both of these goals. That is, it should be able to replicate its original performance on an earlier task when revisiting it after having learned other tasks, all while not interfering with its ability to gain performance on new tasks.
Neural networks have proved a powerful tool driving the state-of-the-art in games, control and interactive language models Vinyals et al. (2019); OpenAI et al. (2019); Beattie et al. (2016); Andrychowicz et al. (2020); Brown et al. (2020). However, neural networks have been shown to be highly susceptible to forgetting previously learnt knowledge after training on a new task French (1999). This undesirable property of neural networks has proven troublesome in their integration into CL systems, since catastrophe forgetting is the most reported failure mode McCloskey and Cohen (1989); Khetarpal et al. (2022). Catastrophic forgetting occurs when newly acquired skills compromise already learned ones. The opposite of catastrophic forgetting is forward and backward transfer, in which either earlier or later tasks benefit from learning on separate tasks.
Typically, the traditional RL settings focus on solving one individual task that does not change during learning Khetarpal et al. (2022). Where multiple different tasks are required to be learned, one common approach is to learn all simultaneously, rather than sequentially. This avoids catastrophic forgetting as the data distribution of the learning task never changes, but it is not always applicable, as the environments must generate too many samples or because some tasks are unavailable until others have completed. This is the case in continual RL, where new challenges are often experienced one after another. Therefore, a CL system must be able to learn multiple different tasks sequentially. Additionally, tasks may change subtly or silently over time, so a CL system should also be capable without explicit task identifiers Rolnick et al. (2019).
One method that existing CL systems employ in RL is to store experiences of all tasks, so that the system may ‘remember’ and retain knowledge from previous tasks to prevent catastrophic forgetting Lipton et al. (2018). This is done through a replay buffer, often implemented as a generic first-in-first-out (FIFO). Given that the buffer is sufficiently large and that the learning system is capable of using the knowledge stored in the buffer (i.e., the RL algorithm is not a purely on-policy method), such a solution may enable favourable properties of CL such as reduced forgetting and forward and backward transfer. However, the requirement to store all experiences from all tasks, demands very high storage capacity. This is an unfavourable property, as RL is often done at scale and the large volume of samples to store can present a significant engineering challenge, reducing the scalability and accessibility of the algorithm OpenAI et al. (2019).
@@
The idea of world models has been exploited in RL algorithms (Model Based RL), where a world model is capable of predicting the effects of actions on the environment Ha and Schmidhuber (2018); Hafner et al. (2019, 2020, 2022, 2023). RL world model approaches are an intuitive choice for CL, as they naturally support off-policy learning; retaining the world model’s knowledge of a task is likely sufficient for preserving performance. Although CLS describes replay to a ‘world model’, traditionally replay in RL is used in Model Free RL to improve the policy, rather than a world model directly.
In this paper, we take the most recent development of world model approaches in DreamerV3 Hafner et al. (2023) and explore its ability in CL when used with memory-efficient replay buffers Isele and Cosgun (2018). Hence, we present World Models with Augmented Replay (WMAR) as a promising solution to CL problems faced by world model approaches, while alleviating memory requirements. We evaluate WMAR on a diverse and challenging set of environments from OpenAI Procgen (procedurally generated games) and Atari. These suites pose distinct challenges as to how similar or dissimilar different tasks are. We used Procgen to create similar tasks with ‘shared structure’, whereas Atari games comprise dissimilar tasks ‘without shared structure’. Our analysis is not restricted to catastrophic forgetting but includes forgetting and backward/forward transfer of skills between tasks.
@@
For a background in RL and CL, including desiderata in CL solutions, see Section S.LABEL:S-sec:background. For Related Work, see Section S.LABEL:S-sec:related_work; including methods to counteract catastrophic forgetting in CL Kirkpatrick et al. (2017); Schwarz et al. (2018); Shin et al. (2017) and continual RL Isele and Cosgun (2018); Rolnick et al. (2019); Kessler et al. (2022).
@@ 2 World Models with Augmented Replay (WMAR)
WMAR uses world models and augmented replay buffers to learn tasks sequentially. The world model builds on developments in DreamerV3, and is state-of-the-art in single-GPU RL benchmarks in a number of tasks. The off-policy nature of the RL algorithm allows for the use of replay buffers as a mechanism to prevent forgetting. Furthermore, world models naturally support a memory-based continual learning approach as the strength of the final policy is largely contingent on the quality of the world model’s accuracy in simulating the environment. Hence, preventing forgetting within the world model through any means should retain agent performance in past environments. World models are also task-agnostic and labels specifying the task/environment are not required.
The system consists of the World Model for modelling the environment, an Actor-Critic Controller to act on the environment, and a Replay Buffer. The Replay Buffer is used to train the World Model, and the World Model is able to ‘dream’ experiences, which are used to train the Controller. Each is described below, and more details are found in Section S.LABEL:S-sec:world_models.
@@ 2.1 World Model
Following DreamerV3, the stochastic state $z_{t}$ consists of 32 discrete stochastic units with 32 categorical classes each. Should the RSSM be able to accurately predict the environment dynamics (i.e., $\hat{z}_{t}\approx z_{t}$), it follows that the deterministic and stochastic states together form a Markovian representation of the undisclosed environment state. During open-loop prediction, stochastic state posteriors $z_{t}$ do not exist and the priors are taken for the posteriors instead (i.e. $z_{t}=\hat{z}_{t}$). We use the standard implementation of the GRU with Tanh activation, as we did not observe a significant improvement at the cost of extended wall time.
@@ 2.2 Actor Critic
The actor and critic (Figure S.LABEL:S-fig:actor_critic) are MLPs mapping the model state to actions and value estimates. They are trained entirely on trajectories generated stochastically by the world model, which is referred to as ‘dreaming’ or ‘imagining’. Starting from model states of inputs sampled from the replay buffer, the actor acts within states predicted by the world model, producing a sequence of states, actions, rewards, and continuation flags. We follow DreamerV3 in training the actor and critic on-policy on these imagined trajectories using reinforcement Williams (1992). As such, inevitable changes in distribution within the model state space are not an issue since imagined trajectories are cheap to generate and do not require interaction with the actual environment, allowing this process to be run to convergence of the actor and critic.
@@ 2.3 Replay Buffer
A key objective of WMAR is to limit the size of the replay buffers. While Dreamer maintained a single FIFO buffer containing the last $10^{6}$ observations, we empirically choose a size of $2^{18}\approx 262$K observations for each buffer, as it is significantly (about 4×4$\times$4 ×) smaller than that of Dreamer’s, without noticeably costing performance. We also introduce spliced rollouts, which are a simple and sometimes necessary alternative to storing entire episodes, enabling smaller buffer sizes. The buffer types are shown in Figure 2 and described below.
@@ 2.3.1 Short-Term FIFO Buffer
We maintain a FIFO buffer with a capacity of $2^{18}$ samples containing the most recent rollout observations. The retention of a short-term FIFO buffer follows CLS theory, suggesting that multiple separate structures within a system are required for learning and CL. This buffer also allows the world model to train on all incoming experiences and biases the training samples to those more recently collected, thus improving convergence properties on the current task.
@@ 2.3.2 Long-Term Global Distribution Matching (LTDM)
Matching the global training distribution even with limited capacity in the replay buffer has been shown to be a suitable method to reduce catastrophic forgetting Isele and Cosgun (2018). We use a long-term global distribution matching buffer, also with a capacity of $2^{18}$ samples that contain a uniformly random subset of 512 spliced rollouts. This buffer uses the reservoir sampling strategy by assigning a random value for each rollout chunk as a key in a size-limited priority queue, preserving the experiences with the highest key values, and discarding the remainder.
@@
Restricting the size of the replay buffer can cause it to become full with only a few episodes, especially as episode lengths may be in the thousands. Containing too few unique episodes may cause the training data to become highly unrepresentative of the general environment states, thus reducing the world model’s accuracy in modelling the environment and likely causing a subsequent loss of performance from the actor. This is especially relevant for the long-term global distribution matching buffer, which may store relatively few samples from each environment.
Hence, while typical implementations store complete rollouts, we splice rollouts into smaller chunks of size 512 where necessary, which we find to not adversely affect performance. The implication is that the long-term global distribution matching buffer will operate over the spliced rollouts rather than entire rollouts, thus providing a guarantee on the granularity for its sampling of data.
Remainders of rollouts with fewer than 512 states are concatenated prior to the next episode with an appropriate reset flag indicating the start of a new episode. To improve the efficiency of the rollout process, we truncate episodes after a fixed number of steps, so that the final number of environment steps is identical in every training iteration. We find spliced rollouts to be a simple solution to control granularity in the distribution matching buffer, especially when episodes are extremely long.
@@ 2.3.3 Augmented Replay Buffer
We combine the equally sized short-term FIFO buffer $\mathcal{D}_{1}$ with the long-term global distribution matching buffer $\mathcal{D}_{2}$ by using both buffers in parallel. Data from both buffers are uniformly sampled for each training minibatch (Algorithm S.LABEL:S-alg:combined_buffers).
@@ 2.4 Task Agnostic Exploration
The exploration of new environments may pose a challenge, as policies that have already been trained on previous tasks may lack the randomness required to adequately explore the state space of a new task. To address this challenge, we use developments from the DreamerV3 algorithm, including the introduction of a world model regulariser and scaling of returns in the actor critic training step. We find that these developments enable mitigating the exploration challenge with a fixed entropy regulariser and are able to produce strong results without the addition of an exploration-oriented learning system such as Plan2Explore (Section S.LABEL:S-sec:task_agnostic). Avoiding the need to train and run inference on an ensemble of predictors also provides benefits in wall time.
@@ 3 Experiments
We perform an empirical study to evaluate the efficacy of WMAR for CL in a set of challenging environments from OpenAI Procgen and Atari Cobbe et al. (2020); Bellemare et al. (2013), which we select for the unique challenges they present in a CL setting, where environments may be extremely similar (Procgen) or dissimilar (Atari). As measures of baseline performance, we compare the performance of WMAR with the most similar model-free and model-based RL algorithms on individual tasks. We choose proximal policy optimisation (PPO) Schulman et al. (2017); Raffin et al. (2021) for the model-free RL algorithm, and use the CNN-LSTM network architecture as it most similarly resembles the architecture of WMAR. We use DreamerV3 as the model-based benchmark.
To measure the efficacy of the long-term distribution matching buffer, we compare CL metrics for WMAR and WMAR with only the FIFO buffer and compare them both with DreamerV3 on Atari. Finally, we compare the performance difference of WMAR when applied to multiple sequential tasks versus when applied to only single tasks. All WMAR experiments are trained on one Nvidia A40 40GB GPU each with all single task benchmarks running within 0.25 days and continual learning benchmarks running within 1 day of wall time, making these experiments widely available and reproducible across research labs.
For more details on experimental parameters and execution time, see Section S.LABEL:S-sec:extime.
@@ 3.1 Evaluation
We evaluate the performance by normalising episodic reward to the performance provided by the world model agent’s baseline, and that of a random agent. This enables a normalised comparison in order to measure the relative performance during CL. These results also enable a natural evaluation of forgetting and forward transfer for all environments in each CL suite.

### 05 Mar 2024 v2

@@ Abstract
Continual RL is a challenging problem where the agent is exposed to a sequence of tasks; it should learn new tasks without forgetting old ones, and learning the new task should improve performance on previous and future tasks. The most common approaches use model-free RL algorithms as a base, and replay buffers have been used to overcome catastrophic forgetting. However, the buffers are often very large making scalability difficult. Also, the concept of replay comes from biological inspiration, where evidence suggests that replay is applied to a world model, which implies model-based RL – and model-based RL should have benefits for continual RL, where it is possible to exploit knowledge independent of the policy. We present WMAR, World Models with Augmented Replay, a model-based RL algorithm with a world model and memory efficient distribution matching replay buffer. It is based on the well-known DreamerV3 algorithm, which has a simple FIFO buffer and was not tested in a continual RL setting. We evaluated WMAR vs WMAR (FIFO only) on tasks with and without shared structure from OpenAI ProcGen and Atari respectively, and without a task oracle. We found that WMAR has favourable properties on continual RL with significantly reduced computational overhead compared to WMAR (FIFO only). WMAR had small benefits over DreamerV3 on tasks with shared structure and substantially better forgetting characteristics on tasks without shared structure; but at the cost of lower plasticity seen in a lower maximum on new tasks. The results suggest that model-based RL using a world model with a memory efficient replay buffer can be an effective and practical approach to continual RL, justifying future work.
@@ 1 Introduction
Typically, RL settings focus on one individual task that does not change during learning. In continual RL, new tasks are experienced one after another, which is more challenging due to the well known issue of catastrophic forgetting McCloskey & Cohen (1989); Khetarpal et al. (2022) but represents many real-world scenarios Khetarpal et al. (2022). A simple approach to learning multiple tasks is to learn them simultaneously, but in continual RL, there are often too many environments for that to be practical and tasks are unavailable until others have completed. Additionally, tasks may change subtly or silently over time, so the system should not require explicit task identifiers Rolnick et al. (2019).
One approach to continual RL is to store experiences of all tasks, so that the system may ‘remember’ and retain knowledge to prevent catastrophic forgetting Lipton et al. (2018). This is often done through a replay buffer, implemented as a generic first-in-first-out (FIFO). However, the requirement to store all experiences from all tasks, demands very high storage capacity, reducing scalability and accessibility of the algorithm OpenAI et al. (2019).
@@
The idea of world models has been exploited in model-based RL algorithms, where a world model is capable of predicting the effects of actions on the environment Ha & Schmidhuber (2018); Hafner et al. (2019) and the more recent state-of-the-art DreamerV1 to V3 Hafner et al. (2020; 2022; 2023). To our knowledge, world model approaches have not been applied to continual RL. However, they are an intuitive choice, as they naturally support off-policy learning; retaining the world model’s knowledge of a task is likely sufficient for preserving performance. Although CLS describes replay to a ‘world model’, traditionally replay in RL is used in Model Free RL to improve the policy, rather than a world model directly.
In this paper, we take DreamerV3 Hafner et al. (2023), add a memory-efficient replay buffer Isele & Cosgun (2018), and apply it to continual RL. Hence, we present World Models with Augmented Replay (WMAR).
We applied WMAR to two settings. First, where each task has a distinct environment and reward function. This is the most comment setting used in continual RL, and Atari games are often used. In the second setting, there is commonality between tasks, and learned knowledge can be leveraged to perform subsequent tasks. This is referred to as tasks with ‘shared structure’ Khetarpal et al. (2022). Often a video game is used, but the conditions such as movement dynamics or spacing of features in the environment change over time e.g. Riemer et al. (2019). Many potential real-world applications exist within the ‘shared structure’ setting. For example a robot assistant that should be able to acquire new and related tasks within a home like cleaning with a broom and then a mop.
We used Procgen (procedurally generated games) from OpenAI to create similar tasks with ‘shared structure’, and Atari games for dissimilar tasks ‘without shared structure’. Our analysis is not restricted to catastrophic forgetting but includes forgetting and backward/forward transfer of skills between tasks.
@@
For background and related work, see Appendix A; including methods to counteract catastrophic forgetting in CL Kirkpatrick et al. (2017); Schwarz et al. (2018); Shin et al. (2017) and continual RL Isele & Cosgun (2018); Rolnick et al. (2019); Kessler et al. (2022).
1.1 Desiderata
Using a neural network, policy updates arising from the need to learn the current task often prove detrimental to preserving knowledge and performance on previous tasks. On the contrary, humans are able to accumulate knowledge to benefit performance on multiple related tasks simultaneously. Hence, a successful CL agent should be able to exhibit the following qualities Khetarpal et al. (2022); Chen & Liu (2018):
  • Stability The agent avoids forgetting and losing performance on previously learnt tasks when learning new tasks.
  • Backward Transfer The agent can increase its performance on previously learned tasks after training on new, similar tasks.
  • Plasticity The agent, when encountering a new task unlike any of the previous, does not learn it any slower compared to learning it independently.
  • Forward Transfer When encountering a new task that is similar to previously learnt tasks, the CL agent can learn it faster compared to learning it independently.
Another challenge of continual RL is the lack of explicit task boundaries. Realistically, tasks may change unexpectedly, or evolve slowly over time. In these cases, a task oracle is either not available or not applicable. Further, practical considerations regarding CL such as its scalability and the wall time of experiments are a relevant point of interest. Therefore, the CL algorithm should also possess the following properties:
  • Simple Integration of multiple systems supporting CL is a possibility in future research. WMAR proposes a simple and direct method of using replay within the context of RL with learned world models which may be combined easily with other approaches. Furthermore, WMAR is simple to implement and imposes minimal modification to the underlying RL algorithm, DreamerV3 Hafner et al. (2023).
  • Scalable The solution should be able to scale to support CL on a large number of tasks. In order to do so, memory and computational constraints are necessary to ensure the wider accessibility and practicality of the solution.
  • Does not require a task oracle A task oracle provides exact knowledge of which task is currently being trained, which may not be possible in practice. The requirement of a task oracle decreases the applicability of the solution.
1.2 Contributions
The primary contributions include: a) applying ‘world model’ DreamerV3 based architectures to the problem of continual RL for tasks with and without shared structure and b) enhancing the replay buffer of DreamerV3.
@@ 2 World Models with Augmented Replay (WMAR)
WMAR uses world models and augmented replay buffers to learn tasks sequentially. The world model builds on developments in DreamerV3 which is state-of-the-art in single-GPU RL benchmarks in a number of tasks. The off-policy nature of the algorithm allows a replay buffer to prevent forgetting. Furthermore, world models naturally support a memory-based continual learning approach as the strength of the final policy is largely contingent on the quality of the world model’s accuracy in simulating the environment. Hence, preventing forgetting within the world model through any means should retain agent performance in past environments. World models are also task-agnostic and labels specifying the task/environment are not required.
The system consists of the World Model for modelling the environment, an Actor-Critic Controller to act on the environment, and a Replay Buffer. The Replay Buffer is used to train the World Model, and the World Model is able to ‘dream’ experiences, which are used to train the Controller. Each is described below, and more details are found in Appendix B.
@@ 2.1 World Model
Following DreamerV3, the stochastic state $z_{t}$ consists of 32 discrete stochastic units with 32 categorical classes each. Should the RSSM be able to accurately predict the environment dynamics (i.e., $\hat{z}_{t}\approx z_{t}$), it follows that the deterministic and stochastic states together form a Markovian representation of the hidden environment state. During open-loop prediction, stochastic state posteriors $z_{t}$ do not exist, and the priors are taken for the posteriors instead (i.e. $z_{t}=\hat{z}_{t}$). We used the standard implementation of the GRU with Tanh activation, as we did not observe a significant improvement at the cost of extended wall time.
@@ 2.2 Actor Critic
The actor and critic (Figure 5, Appendix B.2) are MLPs that map the state of the model to actions and value estimates. They are trained entirely on trajectories generated stochastically by the world model, which is referred to as ‘dreaming’ or ‘imagining’. Starting from model states of inputs sampled from the replay buffer, the actor acts within states predicted by the world model, producing a sequence of states, actions, rewards, and continuation flags. We followed DreamerV3 in training the actor and critic on-policy on these imagined trajectories using reinforcement Williams (1992). As such, inevitable changes in distribution within the model state space are not an issue since imagined trajectories are cheap to generate and do not require interaction with the actual environment, allowing this process to be run to convergence of the actor and critic.
@@ 2.3 Replay Buffer
A key objective of WMAR is to limit the size of the replay buffers. While Dreamer maintained a single FIFO buffer containing the last $10^{6}$ observations, we empirically chose a size of $2^{18}\approx 262$K observations for each buffer, as it is significantly (about 4×4$\times$4 ×) smaller than that of Dreamer’s, without noticeably costing performance. We also introduced spliced rollouts, which are a simple and sometimes necessary alternative to storing entire episodes, enabling smaller buffer sizes. The buffer types are shown in Figure 2 and described below.
@@ 2.3.1 Short-Term FIFO Buffer
A FIFO buffer with a capacity of $2^{18}$ samples contained the most recent rollout observations. The retention of a short-term FIFO buffer follows CLS theory, suggesting that multiple separate structures within a system are required for learning and CL. This buffer also allowed the world model to train on all incoming experiences and biased the training samples to those more recently collected, thus improving convergence properties on the current task.
@@ 2.3.2 Long-Term Global Distribution Matching (LTDM)
Matching the global training distribution even with limited capacity in the replay buffer has been shown to be a suitable method to reduce catastrophic forgetting Isele & Cosgun (2018). We used a long-term global distribution matching buffer, also with a capacity of $2^{18}$ samples that contained a uniformly random subset of 512 spliced rollouts. This buffer used the reservoir sampling strategy by assigning a random value for each rollout chunk as a key in a size-limited priority queue, preserving experiences with the highest key values, and discarding the remainder.
@@
Restricting the size of the replay buffer can cause it to become full with only a few episodes, especially since the duration of the episodes may be in the thousands. Containing too few unique episodes may cause the training data to become highly unrepresentative of the general environment states, thus reducing the world model’s accuracy in modelling the environment and likely causing a subsequent loss of performance from the actor. This is especially relevant for the long-term global distribution matching buffer, which may store relatively few samples from each environment.
Hence, while typical implementations store complete rollouts, we spliced rollouts into smaller chunks of size 512 where necessary, which did not adversely affect performance. The implication is that the long-term global distribution matching buffer operated over the spliced rollouts rather than entire rollouts, thus providing a guarantee on the granularity for its sampling of data.
Remainders of rollouts with fewer than 512 states are concatenated prior to the next episode with an appropriate reset flag indicating the start of a new episode. To improve the efficiency of the rollout process, we truncated episodes after a fixed number of steps, so that the final number of environment steps was identical in every training iteration. We found spliced rollouts to be a simple solution to control granularity in the distribution matching buffer, especially when episodes are extremely long.
@@ 2.3.3 Augmented Replay Buffer
We combined the equally sized short-term FIFO buffer $\mathcal{D}_{1}$ with the long-term global distribution matching buffer $\mathcal{D}_{2}$ by using both in parallel. Data from both buffers were uniformly sampled for each training minibatch (Algorithm 2).
@@ 2.4 Task Agnostic Exploration
Exploring new environments may pose a challenge, as policies that have already been trained on previous tasks may lack the randomness required to adequately explore the state space of a new task. To address this challenge, we used developments from the DreamerV3 algorithm, including the introduction of a world model regulariser and scaling of returns in the actor-critic training step. These developments enabled mitigation of the exploration challenge with a fixed entropy regulariser and produced strong results without the addition of an exploration-orientated learning system such as Plan2Explore (Appendix B.4). Avoiding the need to train and run inference on an ensemble of predictors also benefits wall time.
@@ 3 Experiments
We conducted an empirical study to evaluate the efficacy of WMAR for continual RL in a set of challenging OpenAI Procgen and Atari environments Cobbe et al. (2020); Bellemare et al. (2013). We selected them because they are commonly used as RL benchmarks, can be extremely similar for the case of shared structure (Procgen) or dissimilar in the case of without shared structure (Atari), and they were computationally feasible on the available hardware.
We compared WMAR (with augmented replay buffer) with a baseline model equivalent to DreamerV3, referred to as ‘WMAR (FIFO Only)’. The comparison measures the efficacy of the long-term distribution matching buffer. To validate our implementation, we compared also to an open source implementation of DreamerV3 by the author https://github.com/danijar/dreamerv3, see Appendix D.1.
To measure baseline performance, used for normalisation of other results, we also ran experiments on individual tasks, and compared WMAR with the most similar model-free and model-based RL algorithms. We chose proximal policy optimisation (PPO) Schulman et al. (2017); Raffin et al. (2021) for the model-free RL algorithm and used the CNN-LSTM network architecture, as it most closely resembles the WMAR architecture.
Experimental setup
All WMAR experiments are trained on one Nvidia A40 40GB GPU with all single task benchmarks running within 0.25 days and continual learning benchmarks running within 1 day of wall time, making these experiments widely available and reproducible across research labs.
For more details on experimental parameters and execution time, see Appendix C.2.
@@ 3.1 Evaluation
We evaluated the performance by normalising episodic reward to the performance provided by the world model agent’s baseline and that of a random agent. This enables a normalised comparison in order to measure the relative performance during CL. These results also enable a natural evaluation of forgetting and forward transfer for all environments in each CL suite. More details are given in the following subsections.